The perimeter of a rectangle is 150 cm. The rectangle width is half of its lenght. How long is the width?
Answer by yogshweta(2)

Let length is x cm
Then width is x/2
Perimeter is 150 cm
Perimeter =2*(length +width)
150=2*(x+x/2)
150=2*(3x/2)
150=3x
x=50
Ans :width is x/2=50/2=25 cm